# Pathways of Thoughts: Multi-Directional Thinking for Long-form Personalized Question Answering

### v1

@@ 4.2. Multi-Directional Thinking by Diverse Planning and Mixture-of-Responses
Table 1. Performance on LaMP-QA. $\dagger$ and $\ddagger$ show statistically significant improvement over the single and multiple inference best baseline, respectively, using t-test ($p<0.05$). In this experiment, Gemini 1.5 Pro has been used as the backbone LLM for all the methods.
Method Arts & Lifestyle & Personal Society & Average Entertainment Development Culture (macro) Single Inference (1) No Personalization 0.1741 0.2863 0.2996 0.2533 (2) In-Context Personalization 0.1722 0.3022 0.3482 0.2742 (3) In-Context Personalization w/ CoT 0.2119 0.3862 0.4230 0.3403 (4) Pathways of Thoughts ($N=1$, $T=8$) 0.2499 0.3966 0.4490 0.3651 Multiple Inference (5) In-Context Personalization w/ Best-of-32 0.2248 0.3753 0.4303 0.3434 (6) In-Context Personalization w/ CoT & Best-of-32 0.2607 0.4380 0.4789 0.3925 (7) Tree of Thoughts ($K=32$, $B=2$) 0.2054 0.4178 0.4546 0.3592 (8) Pathways of Thoughts ($N=16$, $T=8$) 0.2999†‡ 0.4960†‡ 0.5362†‡ 0.4440†‡
@@
  • Initial State Alteration: If the initial state is modified before executing the planning action, the LLM can produce different plans, as the changes in personalized context influences the planning process. This can be modeled by varying the amount of past user history accessible to the agent, effectively addressing the challenge of long personal context. To implement this, we randomly select $N$ subsets of the user profile $P_{u}$, each of size $\tau\times|P_{u}|$, where $0\leq\tau\leq 1$. Starting from these $N$ diverse initial states lead to different planning trajectories and thinking pathways.
@@
  • Mixture-of-N: Given the extracted important aspects $I_{u}$ for the user $u$ and the input $x_{u}$, the LLM $\pi$ combines different aspects of the generated responses from the set of multi-directional thinking outputs $R_{\text{PoT}}$. The final response is denoted as $\hat{y}_{x_{u}}=\pi(x_{u},R_{\text{PoT}},I_{u})$, using the prompt shown in Figure 10 in Appendix A.2. This mimics how humans solve problems by exploring multiple solutions and synthesizing a new solution that incorporates the strengths of each solution. This is particularly beneficial for open-ended text generation, where there is no definitive right or wrong response. By aggregating the strengths of all responses, it generates a response that aligns with user preferences, ensuring a more personalized output. In PoT, we use this approach as the primary response aggregation method.
Figure 3. Effect of the number of pathways on the performance.
@@ 5.1. Experimental Setup
We conduct experiments on the LaMP-QA benchmark (Salemi and Zamani, 2025a), the only publicly available dataset for personalized QA. LaMP-QA covers three diverse domains: (1) Art & Entertainment (767 questions), (2) Lifestyle & Personal Development (989 questions), and (3) Society & Culture (1074 questions). [^5] Other datasets for personalized text generation, such as LaMP (Salemi et al., 2024b) and LongLaMP (Kumar et al., 2024), are not suitable for our setting for two main reasons. First, they primarily target open-ended content generation, which differs from the QA focus of this work (Section 1). Second, unlike LaMP-QA, which provides explicit ground-truth annotations of user expectations, they include only a single reference response. This restricts their ability to capture the diversity of valid personalized outputs and can lead to misalignment with user expectations, as noted by Salemi and Zamani (2025a). These limitations are particularly problematic for our feedback-based approach, where convergence to a single reference response is both infeasible and suboptimal for reliable evaluation. Each LaMP-QA example contains a user query, the user’s question history (serving as a profile), a narrative that reflects the user’s perspective and intent, and a set of personalized rubrics specifying the aspects an ideal response should address. For evaluation, we adopt the procedure of Salemi and Zamani (2025a), but replace their suggested evaluator (Qwen-2.5, 32B) with Gemini 1.5 Pro, a more capable model. For each question, the evaluator checks whether the generated response addresses each personalized aspect, assigning scores in the range $[0,2]$. These are normalized to $[0,1]$, and the final response score is computed as the mean normalized score across all aspects. For details, we refer the reader to Salemi and Zamani (2025a). In one experiment, we corroborate our findings through a side-by-side human evaluation comparing PoT to the strongest baseline.
@@
We use Gemini 1.5 Pro as the backbone LLM. To show the generalizability of PoT, we use GPT-4o-mini [^6] Available at: https://platform.openai.com/docs/models/gpt-4o-mini (OpenAI, 2024) using our most representative experiment configurations. For single pathways, we use Nucleus sampling (Holtzman et al., 2020) with a temperature of $\tau=0.1$. For multi-directional pathways, we use the same setting, except for the Planning Action Variation, where a temperature of $\tau=0.9$ is used exclusively for the planning action. For generating diverse plans, we use Planning Action Variation by default and Mixture-of-N for response aggregation, unless otherwise specified. We set the context length to $32k$, the maximum number of actions per pathway to $T=8$ , and the number of pathways to $N=16$. In ablations, we reduce the number of pathways to $N=4$ to save costs, unless otherwise specified. For personalized context, we use the 10 recent questions of the user with the same order as they appear in the LaMP-QA (Salemi and Zamani, 2025a) dataset.
@@
We use 6 baselines: (1) LLM without personalized context to directly answers the question, (2) LLM with access to personalized context, and (3) LLM with access to personalized context that first performs reasoning over the context to extract important aspects relevant to the question before generating the response. The two latter use the prompts shown in Figures 12 and 13 in Appendix B. These baselines all use Nucleus sampling with a temperature of $\tau=0.1$. We extend the latter two baselines by doing multiple inferences and Best-of-N (Ichihara et al., 2025; Brown et al., 2024) to evaluate their performance when test-time compute scales, to form baselines (4) and (5). Additionally, we include the (6) Tree of Thoughts (ToT) (Yao et al., 2023) as a baseline, leveraging its test-time compute scaling. We follow their setup, particularly in the creative writing task, as it is the closest task they studied to personalized question answering. This setup consists of an intermediate planning step, where a set of plans is generated and the best plan is selected, followed by a final response generation step, where a set of responses is generated and the best response is chosen. For planning, we use the prompts shown in Figure 14, and for response generation, we refer to Figure 15, both in Appendix B. We observed that the baselines generate, on average, half the number of tokens compared to PoT{PoT}. To ensure a fair comparison, we set the best of N baselines to generate double the number of responses, with $N=32$ and a sampling temperature of $\tau=0.9$ to ensure diverse responses. For the Best-of-N approach, we utilize the method introduced in Section 4.2. For ToT, following Yao et al. (2023) and their setup for the creative writing task that is the most similar to our task, we set the tree depth to $B=2$. For fair comparison with PoT, we generate $N=32$ plans and responses at each node with a sampling temperature of $\tau=0.9$. The prompts for selecting the best plan is shown in Figure 16 in Appendix B. For selecting the best response, we utilize the method introduced in Section 4.2.
@@ 5.2. Empirical Findings
The results of this experiment are reported in Table 1. The findings show that PoT statistically significantly outperforms all the baselines across all categories, achieving a 13.1% relative improvement on average over the best baseline (row 6). This clearly demonstrates the superiority of PoT compared to the baselines. To compare our thinking-as-an-MDP method with baselines, we also report the results using only a single inference (rows 1-4). The results from this experiment show that our method (row 4) outperforms all single inference baselines (rows 1-3), with statistically significant improvements in two out of three categories and overall performance. Specifically, our method achieves a 7.2% relative improvement in average performance, suggesting that our thinking approach is significantly more effective than CoT.
@@
For this, we focus on a single inference ($N=1$) using PoT and vary the maximum number of actions it can perform. The performance across all categories is shown in Figure 2. The findings show that increasing the number of actions improves performance up to a certain point (8 actions per pathway), after which performance declines. This occurs because, with too few actions, the LLM cannot construct effective pathways to solve the problem. Conversely, when given too many actions, the LLM may overthink the problem, performing unnecessary actions that degrade performance.
@@
We set the maximum number of actions per pathway to $T=8$ and vary the number of pathways. The performance across all categories is shown in Figure 3. The findings indicate that increasing the number of pathways consistently improves performance; however, as the number grows, the improvements become smaller. This occurs because there is a limited set of different ways for answering the question, and after a certain point, the solutions start to repeat, leading to diminishing returns in performance with additional pathways.
@@
We randomly sample 100 examples from LaMP-QA and generate outputs using PoT and In-Context Personalization w/ CoT & Best-of-32. Human annotators evaluate the outputs based on their alignment with the the user’s question narrative. Each example is annotated twice, yielding an inter-annotator agreement Cohen’s kappa of $0.613$. The results indicate that annotators preferred PoT in 66% of the cases, the baseline in 19%, and reported tie in 15% of cases. These results show that PoT achieves considerably higher alignment with human preferences.
@@ 5.3. Case Study
We observe that PoT ($N=16$, $T=8$) generates $1943$ unique pathways, defined by distinct action sequences chosen by the LLM. Specifically, Art & Entertainment yields $1109$, Lifestyle & Personal Development produces $814$, and Society & Culture results in $1182$ unique pathways. On average, each example generates $11.1$ unique pathways with on average $7.19$ actions. As a case study demonstrating the effectiveness of PoT, we present an example in Figure 6:
@@ 6. Conclusions
This paper addresses the critical yet under-explored problem of personalized question answering. we propose Pathways of Thoughts (PoT), a novel approach that models the LLM’s thinking process as a Markov Decision Process (MDP) during test time. By enabling the exploration of multiple thinking pathways and aggregating diverse responses, PoT effectively captures user preferences to generate highly personalized answers. Our experimental results on LaMP-QA demonstrate that PoT achieves substantial improvements over baseline methods, with performance gains of up to 13.1%. Furthermore, human evaluation confirms the effectiveness of our approach, showing a strong user preference over existing baselines. This paper focuses solely on personalized question answering; however Pathways of Thoughts is a general framework and future work can explore the potential of Pathways of Thoughts for long-form text generation tasks beyond personalization.
@@ A.1. Actions Definition
  • checking for personalization: This action asks the model to evaluate whether the user’s question can benefit from personalization. It is useful when determining whether personalizing the response is necessary or if a generic response would suffice.
@@ A.2. Details of Response Aggregation Methods
Figure 14. Prompt for the planning step in ToT.
@@
Appendix B Baselines Prompts
Figures 12 and 13 present the prompts used for the personalized baselines without and with chain-of-thought reasoning, respectively. For the Tree of Thoughts baseline, Figure 14 shows the planning prompt, Figure 15 illustrates the response generation prompt, and Figure 16 provides the plan selection prompt.

### v2

@@ 4.2. Multi-Directional Thinking by Diverse Planning and Mixture-of-Responses
Table 1. Performance on LaMP-QA. $\dagger$ and $\ddagger$ show statistically significant improvement over the single and multiple inference best baseline, respectively, using t-test ($p<0.05$). Gemini 1.5 Pro has been used as the backbone LLM for all the methods.
Method Arts & Lifestyle & Personal Society & Average Entertainment Development Culture (macro) Single Inference (1) No Personalization 0.1741 0.2863 0.2996 0.2533 (2) In-Context Personalization 0.1722 0.3022 0.3482 0.2742 (3) In-Context Personalization w/ CoT 0.2119 0.3862 0.4230 0.3403 (4) Pathways of Thoughts ($N=1$, $T=8$) 0.2499 0.3966 0.4490 0.3651 Multiple Inference (5) In-Context Personalization w/ Best-of-32 0.2248 0.3753 0.4303 0.3434 (6) In-Context Personalization w/ CoT & Best-of-32 0.2607 0.4380 0.4789 0.3925 (7) In-Context Personalization w/ Mix-of-32 0.2310 0.3781 0.4411 0.3500 (8) In-Context Personalization w/ CoT & Mix-of-32 0.2689 0.4443 0.4891 0.4007 (9) Tree of Thoughts ($K=32$, $B=2$) 0.2054 0.4178 0.4546 0.3592 (10) Pathways of Thoughts ($N=16$, $T=8$) 0.2999†‡ 0.4960†‡ 0.5362†‡ 0.4440†‡
@@
  • Initial State Alteration: If the initial state is modified before executing the planning action, the LLM can produce different plans, as the changes in personalized context influences the planning process. This can be modeled by varying the amount of past user interactions accessible to the agent, effectively addressing the challenge of long personal context. To implement this, we randomly select $N$ subsets of the user profile $P_{u}$, each of size $\tau\times|P_{u}|$, where $0\leq\tau\leq 1$. Starting from these $N$ diverse initial states lead to different planning trajectories and thinking pathways.
@@
  • Mixture-of-N: Given the extracted aspects $I_{u}$ for the user $u$ and the input $x_{u}$, the LLM $\pi$ combines different aspects of the generated responses from the set of multi-directional thinking outputs $R_{\text{PoT}}$. The final response is denoted as $\hat{y}_{x_{u}}=\pi(x_{u},R_{\text{PoT}},I_{u})$, using the prompt shown in Figure 10 in Appendix A.2. This mimics how humans solve problems by exploring multiple solutions and synthesizing a new solution that incorporates the strengths of them. This is beneficial for open-ended text generation, where there is no definitive right or wrong response. By aggregating the strengths of all responses, it generates a response that aligns with user preferences, ensuring a more personalized output. PoT uses this approach as the primary response aggregation method.
@@ 5.1. Experimental Setup
We conduct experiments on the LaMP-QA benchmark (Salemi and Zamani, 2025a), the only publicly available dataset for long-form personalized QA. LaMP-QA covers three diverse domains: (1) Art & Entertainment (767 questions), (2) Lifestyle & Personal Development (989 questions), and (3) Society & Culture (1074 questions). Each LaMP-QA example contains a user question, the user’s question history (serving as the user profile), a user-written narrative that reflects the user’s information need and intent, and a set of personalized rubrics specifying the aspects an ideal response should address. For evaluation, we follow the evaluation recipe of Salemi and Zamani (2025a) with Gemini 1.5 Pro as the evaluator language model. For each question, the evaluator checks whether the generated response addresses each personalized aspect, assigning scores in the range $[0,2]$. These are normalized to $[0,1]$, and the final response score is computed as the mean normalized score across all aspects. For details, we refer the reader to Salemi and Zamani (2025a). In one experiment, we corroborate our findings through a side-by-side human evaluation comparing PoT to the strongest baseline.
@@
We use Gemini 1.5 Pro as the backbone LLM. To show the generalizability of PoT, we also use GPT-4o-mini [^2] Available at: https://platform.openai.com/docs/models/gpt-4o-mini (OpenAI, 2024a) using our most representative experiment configurations. For single pathways, we use Nucleus sampling (Holtzman et al., 2020) with a temperature of $\tau=0.1$. For multi-directional pathways, we use the same setting, except for the Planning Action Variation, where a temperature of $\tau=0.9$ is used exclusively for the planning action. For generating diverse plans, we use Planning Action Variation by default and Mixture-of-N for response aggregation, unless otherwise specified. We set the context length to $32k$, the maximum number of actions per pathway to $T=8$ , and the number of pathways to $N=16$. In ablations, we reduce the number of pathways to $N=4$ to save costs, unless otherwise specified. For personalized context, we use the 10 recent questions of the user with the same order as they appear in the LaMP-QA (Salemi and Zamani, 2025a) dataset.
Figure 4. Effect of the number of pathways on the performance.
@@
We use six baselines: (1) LLM without personalized context that directly answers the question, (2) LLM with access to the same personalized context as PoT, and (3) LLM with access to personalized context that first performs reasoning over the context to extract important aspects relevant to the question before generating the response. The two latter use the prompts shown in Figures 11 and 12 in Appendix B. These baselines all use Nucleus sampling with a temperature of $\tau=0.1$. We extend the latter two baselines by doing multiple inferences and Best-of-N (Ichihara et al., 2025; Brown et al., 2024) and Mixture-of-N to evaluate their performance when test-time compute scales, to form baselines (5), (6), (7), and (8). Additionally, we include the (9) Tree of Thoughts (ToT) (Yao et al., 2023) as a baseline, leveraging its test-time compute scaling for comparison with our method. We follow their setup, particularly in the creative writing task, as it is the closest task they studied to personalized question answering. This setup consists of an intermediate planning step, where a set of plans is generated and the best plan is selected, followed by a final response generation step, where a set of responses is generated and the best response is chosen. For planning, we use the prompts shown in Figure 13, and for response generation, we refer to Figure 14, both in Appendix B. We observed that the baselines generate, on average, half the number of tokens compared to PoT{PoT}. To ensure a fair comparison, we set the Best-of-N and Mixture-of-N baselines to generate double the number of responses, with $N=32$ and a sampling temperature of $\tau=0.9$ to ensure diverse responses. For the Best-of-N and Mixture-of-N approach, we use the method introduced in Section 4.2. For ToT, following Yao et al. (2023) and their setup for the creative writing task that is the most similar to ours, we set the tree depth to $B=2$. For fair comparison with PoT, we generate $N=32$ plans and responses at each node with a sampling temperature of $\tau=0.9$. The prompts for selecting the best plan is shown in Figure 15 in Appendix B. For selecting the best response, we utilize the method introduced in Section 4.2.
@@ 5.2. Empirical Findings
The results of this experiment are reported in Table 1. The findings show that PoT outperforms all the baselines across all categories, achieving a 10.8% relative improvement on average over the best baseline (row 8). The improvements are statistically significant with respect to t-test with 95% confidence. This demonstrates the superiority of PoT compared to the baselines. To compare our thinking-as-an-MDP method with baselines, we also report the results using only a single inference (rows 1-4). The results from this experiment show that our method (row 4) outperforms all single inference baselines (rows 1-3), with statistically significant improvements in two out of three categories and overall performance. Specifically, our method achieves a 7.2% relative improvement in average performance, suggesting that our thinking approach is significantly more effective than CoT.
@@
For this, we focus on a single inference ($N=1$) using PoT and vary the maximum number of actions it can perform. The performance across all categories is shown in Figure 3. The findings show that increasing the number of actions improves performance up to a certain point (8 actions per pathway), after which performance declines. This occurs because, with too few actions, the LLM cannot construct effective pathways to solve the problem. Conversely, when given too many actions, the LLM may overthink the problem, performing unnecessary actions that degrade performance (Aggarwal et al., 2025).
@@
We set the maximum number of actions per pathway to $T=8$ and vary the number of pathways. The performance across all categories is shown in Figure 4. The findings indicate that increasing the number of pathways consistently improves performance; however, as the number grows, the improvements become thinner. This occurs because there is a limited set of different ways for answering the question, and after a certain point, the solutions start to repeat, leading to diminishing returns in performance with additional pathways.
@@
We randomly sample 100 examples from LaMP-QA and generate outputs using PoT and In-Context Personalization w/ CoT & Best-of-32. Human annotators evaluate the outputs based on their alignment with the the user’s question narrative (i.e., a detailed description of user’s information need and perspective). Each example is annotated twice, yielding an inter-annotator agreement Cohen’s kappa of $0.613$. The results indicate that annotators preferred PoT in 66% of the cases, the baseline in 19%, and reported tie in 15% of cases. These results show that PoT achieves considerably higher alignment with human preferences.
@@ 5.3. Case Study
We observe that PoT ($N=16$, $T=8$) generates $1943$ unique pathways, defined by distinct action sequences chosen. Specifically, Art & Entertainment yields $1109$, Lifestyle & Personal Development produces $814$, and Society & Culture results in $1182$ unique pathways. On average, each example generates $11.1$ unique pathways with on average $7.19$ actions. As a case study demonstrating the effectiveness of PoT, we present an example in Figure 16 in Appendix:
@@ 6. Conclusions
This paper addresses the critical yet under-explored problem of personalized question answering. we propose Pathways of Thoughts (PoT), a novel approach that models the LLM’s thinking process as an MDP during test time. By enabling the exploration of multiple thinking pathways and aggregating diverse responses, PoT effectively captures user preferences to generate personalized answers. Our results on LaMP-QA demonstrate that PoT achieves substantial improvements over baseline methods, with performance gains of up to 10.8%. Furthermore, human evaluation confirms the effectiveness of our approach, showing a strong user preference over existing baselines. This paper focuses solely on personalized question answering; however Pathways of Thoughts is a general framework and future work can explore the potential of Pathways of Thoughts for long-form text generation beyond personalization.
Acknowledgement
This work was supported in part by the Center for Intelligent Information Retrieval, in part by NSF grant #2143434, in part by the Office of Naval Research contract #N000142412612, and with support from by Google.org. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsor.
@@ A.1. Actions Definition
  • personalization: This action asks the model to evaluate whether the user’s question can benefit from personalization. It is useful when determining whether personalizing the response is necessary or if a generic response would suffice.
@@
A.2. Details of Response Aggregation Methods
@@
Figure 13. Prompt for the planning step in ToT.
Appendix B Baselines Prompts
Figures 11 and 12 present the prompts used for the personalized baselines without and with chain-of-thought reasoning, respectively. For the Tree of Thoughts baseline, Figure 13 shows the planning prompt, Figure 14 illustrates the response generation prompt, and Figure 15 provides the plan selection prompt.